AS/AD Model.
AD: P = 90 - 3Y.
AS: P = 15.
PO: Y = 30.
What is the Recessionary Gap??
1
Expert's answer
2016-05-14T12:37:03-0400
Find macroeconomic equilibrium, when total demand=total supply
AD = AS
P=90-3Y=15
Y (actual GDP) =25
When actual GDP is below the potential GDP the economy has negative output gap, also called as recessionary gap.
Recessionary Gap = Potential GDP-Actual GDP=30-25=5
We received negative gap, also called as recessionary gap.
The recessionary gap equals 5.